Glass Art Manufacturing Berhad has a beta of 1.50, the risk-free rate of interest is currently 12 percent, and the required return on the market portfolio is 18.00 percent. The company plans to pay a dividend of RM1.96 per share in the coming year and anticipates that its future dividends will increase at an annual rate consistent with that experienced over the 2016-2018 period.

Year                      Dividend (RM)

2016                     1.68

2017                     1.77

2018                     1.86

Estimate the value of Glass Art Manufacturing's stock. (Remark: Please workout the calculation with the clear formula)
Answer #1

FV = PV (1+r)^n

Where r is growth rate and n is time gap

1.86 = 1.68(1+r)^2

(1+r)^2 = 1.86/1.68

= 1.1071

1+r = (1.1071)^(1/2)

= 1.0522

r = 1.0522 - 1

= 0.0522 i.e 5.22%

Given D1 = RM 1.96

D2 = D1(1+g)

= RM1.96 (1+0.0522)

= RM 1.96 * 1.0522

= RM 2.0623

Ke = Rf + beta ( Rm - Rf )

Ke = Cost of Equity

Rf = Risk free rate

Rm = Market Ret

Beta = Systematic Risk

Ke = Rf + beta ( Rm - Rf )

= 12% + 1.5 (18% - 12% )

= 12% + 1.5 (6%)

= 12% + 9%

= 21%

P1 = D2 / [ Ke - g ]

D2 = Div after 2 Years

P1 = Price after 1 Year

P1 = RM 2.0623 / [ 21% - 5.22% ]

= RM 2.0623 / 15.78%

= RM 13.07

P0 = [ D1 + P1 ] / [ 1+Ke ]

= [ RM 1.96 + RM 13.07 ] / [ 1 + 0.21 ]

= RM 15.03 / 1.21

= RM 12.42

Pl;s comment, if any further assistance is required.